# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 2.1.1. Attribution Logic
Last-Touch Attribution (LTA), also known as Last-Click Attribution, assigns full credit for a conversion to the last interaction or touchpoint that directly led to the conversion. First-Touch Attribution (FTA), as the name suggests, attributes full credit for a conversion to the first interaction or touchpoint that introduced the customer to the brand or campaign. Uniform Attribution (UNI) is an attribution model in advertising measurement that evenly distributes credit for conversions across all touchpoints in a customer’s journey, providing equal weight to each interaction.
Table 5 shows an example of different attribution logics, given the joined Table 4. Since only ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ is relevant to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$, both LTA, FTA, and UNI will assign a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ (similarly for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$). In the case of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$, where both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ are relevant, LTA attributes a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ as it occurs last. FTA assigns a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ since it happens first. UNI, on the other hand, evenly distributes the weight, assigning 0.5 to both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$.
Because LTA is the most commonly used attribution logic (choi2020online, ), we will use LTA for illustration throughout this paper. However, the techniques discussed can also be applied to other MTA logics, as demonstrated in the experiments (Section 6.3). More sophisticated algorithms, such as incremental value heuristics and Shapley value methods (singal2019shapley, ), will be considered for future work. These algorithms do not focus on a single user’s data but instead consider correlations across the dataset, making their underlying models significantly different from our current settings.
Table 5. Attribution on Day 1 in Table 4
@@
2.1.2. Measuring Attributed Conversions
@@
With LTA, each conversion is attributed to the impression that occurs last. In order to effectively evaluate the advertisement’s performance across multiple publishers, it is crucial to aggregate the total attribution weights, i.e., total conversion, for each publisher. A high cumulative sum of weights for a publisher signifies that advertisements on that platform have a substantial likelihood of resulting in product conversions. This metric becomes instrumental in assessing the effectiveness of ad campaigns and optimizing advertising strategies.
@@
For a specific publisher and conversion event ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$, the attribution process operates as follows: If ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$ is attributed to an impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$, and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ originates from the same publisher (${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ occurs within the publisher’s domain), the publisher is assigned an attribution weight of $w=1$. Conversely, if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ is associated with a different publisher, the attribution weight assigned to the original publisher is $w=0$. Then we will compute, for all conversions, the total attribution weights received by each publisher. A high sum of attribution weights for a publisher signifies that advertisements on that platform are more likely to be profitable and convert into actual conversions.
2.1.3. Streaming Measurement
Table 6. Total Conversion via LTA
Pub-ID Day 1 Day 1 to 2 Day 1 to 3 $\cdots$
P-1 $\mathbf{x}_{\text{P1}}[1]$ $\mathbf{x}_{\text{P1}}[1]+\mathbf{x}_{\text{P1}}[2]$ $\sum_{i=1}^{3}\mathbf{x}_{\text{P1}}[i]$cdots
P-2 $\mathbf{x}_{\text{P2}}[1]$ $\mathbf{x}_{\text{P2}}[1]+\mathbf{x}_{\text{P2}}[2]$ $\sum_{i=1}^{3}\mathbf{x}_{\text{P2}}[i]$cdots
In practical scenarios, it is crucial to implement a continuous approach to releasing the sum of attribution weights. This allows for ongoing monitoring of trends over time, enabling the assessment of changes in advertising performance. Table 6 illustrates the cumulative attribution weight, i.e., total conversion, of Nike across two major publishers, Publisher 1 and Publisher 2. This dataset enables the advertiser, Nike, to assess the performance of their advertisements, a crucial factor directly influencing profitability. The continuous online release of attribution data facilitates real-time insights, thereby empowering businesses to make timely and informed decisions that can significantly impact their bottom line.
Table 6 is an example of a private dataset requiring protection. There are $n$ days in total, and we represent the data vector as $\mathbf{x}\in\mathbb{R}^{n}$, where $\mathbf{x}[i]$ represents the sum of attribution weights for a specific publisher on day $i$ ($1\leq i\leq n$). In particular, the sum of attribution weights is equal to the number of conversions if using Last Touch Attribution or First Time Attribution.
@@ 2.1.3. Streaming Measurement
For each publisher, we aim to calculate the sum of attribution weights from day 1 to day $i$, offering insights into advertisement performance over a specific timeframe. In mathematical terms, our goal is to release the prefix-sum of $\mathbf{x}$, i.e., $(\mathbf{x}[1],\mathbf{x}[1]+\mathbf{x}[2],\cdots,\sum_{i=1}^{n}\mathbf{x}[i])$. Our primary concern is to ensure that the released information will be privacy-preserving. Specifically, we require that the released information satisfies differential privacy, which will be discussed next in Section 2.2. In Section 4, we will introduce the methodology for achieving this differential privacy-preserving release of the private dataset.
2.2. Differential Privacy
@@
Differential Privacy (dwork06Calibrating, ; dworkKMM06:ourdata, ; zcdp, ; renyidp, ; fdp, ) establishes a set of constraints governing the operation of a mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$. It has emerged as the de facto standard for safeguarding confidentiality when generating publicly accessible data products. Its adoption extends to a growing number of real-world applications, with notable examples including the U.S. Census Bureau (ashwin08:map, ; onthemap, ; tdahdsr, ), Uber (elasticsensitivity, ; chorus, ), Apple (appledpscale, ), Facebook (fburlshares, ; opacus, ), Microsoft (DingKY17, ), and Google (rappor, ; tensorflowprivacy, ; googlesql, ). Differential privacy offers a robust guarantee of plausible deniability by constraining an adversary’s ability to ascertain whether a specific individual’s data was included in the dataset or not.
@@
One of the key features of differential privacy is its robust privacy guarantee, which remains effective regardless of an adversary’s knowledge or actions when attempting to attack the data. In this context, an adversary refers to any person or entity trying to extract sensitive information about individuals from the results of a data analysis. This privacy guarantee is maintained even if the adversary possesses unlimited computing power and has full knowledge of the algorithm and system used for data collection and analysis.
@@
Definition 2.1 (Differential Privacy (dworkKMM06:ourdata, )).
@@
Definition 2.2 (Zero-Concentrated Differential Privacy (zCDP) (zcdp, )).
@@
Theorem 2.3 (zCDP to DP Conversion (canonne2020discrete, )).
@@
The $L_{2}$ sensitivity of a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is given by,
$\Delta_{2}(f)=\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}}\|f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})-f(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\|_{2},$
where $\|\mathbf{v}\|_{2}$ is the L2 norm of the vector $\mathbf{v}$.
Theorem 2.5 (Gaussian Mechanism of zCDP (zcdp, )).
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be a function of the dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, the Gaussian mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ adds i.i.d. Gaussian noise to $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ with scale parameter $\sigma^{2}$,
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})+N(0,\sigma^{2}\mathbf{I}_{n}),$
where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. The Gaussian mechanism satisfies $\frac{\Delta_{2}(f)^{2}}{2\sigma^{2}}$-zCDP.

### v2

@@ 2.1.1 Attribution Logic
TABLE V: Attribution on Day 1 in Table IV
@@
2.1.2 Measuring Attributed Conversions
@@
In order to effectively evaluate the advertisement’s performance across multiple publishers, it is crucial to aggregate the total attribution weights, i.e., total conversions, for each publisher. A large cumulative sum of weights for a publisher signifies that advertisements on that platform have a substantial likelihood of resulting in product conversions. This metric becomes instrumental in assessing the effectiveness of ad campaigns and optimizing advertising strategies.
@@
For a specific publisher and conversion event ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$, the attribution process operates as follows: If ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$ is attributed to an impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ from this publisher, the publisher is assigned an attribution weight of $w$. Then we will compute, for all conversions, the total attribution weights received by each publisher. A large sum of attribution weights for a publisher signifies that advertisements on that platform are more likely to be profitable and convert into actual conversions.
2.1.3 Streaming Measurement
TABLE VI: Streaming Measurement of Attributed Conversions across Publishers via Queries
Pub-ID Day 1 Day 2 Day 3 $\cdots$
P-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1}(\mathbf{x}_{\text{P1}}[1])$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{2}(\mathbf{x}_{\text{P1}}[1:2])$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{3}(\mathbf{x}_{\text{P1}}[1:3])$cdots
P-2 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1}(\mathbf{x}_{\text{P2}}[1])$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{2}(\mathbf{x}_{\text{P2}}[1:2])$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{3}(\mathbf{x}_{\text{P2}}[1:3])$cdots
In practical applications, it is essential to adopt a continuous approach for releasing queries on attribution weights. For instance, an advertiser might want to obtain the cumulative sum of attribution up to the current day and report this sum daily. Alternatively, they might be interested in reporting the sum of attributions for the past week, using a sliding window approach. This approach facilitates the ongoing monitoring of trends over time, allowing for the evaluation of changes in advertising performance.
Before presenting an example, let’s define some key notations. We assume the ad campaign will run for $n$ days, and we represent the data vector as $\mathbf{x}_{\text{P}}\in\mathbb{R}^{n}$, where $\mathbf{x}_{\text{P}}[i]$ denotes the sum of attribution weights for publisher P on day $i$ ($1\leq i\leq n$). For simplicity, we use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to represent the data vector when the publisher P is clear from the context. On day $i$, the advertiser is interested in obtaining an answer to the query ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1:i])$. The function ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}$ depends on ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1:i]$ and has access only to the first $i$ entries of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, as the remaining entries are yet to be determined.
Table VI provides an example of streaming measurement for attributed conversions using online queries. In this scenario, the advertiser Nike is interested in obtaining answers to online queries ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{2},\dots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{n}$ across two major publishers, Publisher 1 and Publisher 2. This dataset allows Nike to evaluate the performance of their advertisements in real-time, enabling Nike to make timely and informed decisions that can significantly impact their business.
@@ 2.1.3 Streaming Measurement
Our primary concern is to ensure that the released information will be privacy-preserving. Specifically, we require that the released information satisfies differential privacy, a formal privacy guarantee that will be discussed next in Section 2.2. In Section 4, we will introduce the methodology for achieving this privacy-preserving release of the dataset.
2.2 Differential Privacy
@@
Differential Privacy [20, 8] establishes a set of constraints governing the operation of a mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$. It has emerged as the de facto standard for safeguarding data privacy when publicly accessible data are released. Differential privacy offers a robust guarantee of plausible deniability by constraining an adversary’s ability to ascertain whether a specific individual data record was included in the dataset or not.
@@
One of the key features of differential privacy is its robust privacy guarantee, which remains effective regardless of an adversary’s knowledge or actions when attempting to attack the data. In this context, an adversary refers to any person or entity trying to extract sensitive information about an individual record from the analysis results of the data. This privacy guarantee is maintained even if the adversary possesses unlimited computing power and has full knowledge of the algorithm and system used for data collection and analysis.
@@
Definition 2.1 (Differential Privacy [20]).
@@
When $\delta=0$, we say ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $\epsilon$-differential privacy.
@@
Definition 2.2 (Zero-Concentrated Differential Privacy (zCDP) [8]).
@@
The following two theorems show the conversion between zCDP and $(\epsilon,\delta)$-DP, $\epsilon$-DP.
Theorem 2.3 (zCDP to DP Conversion [9]).
@@
Theorem 2.4 (Pure DP to zCDP Conversion [43]).
If mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $\epsilon$-DP, then it satisfies $\rho$-zCDP with $\rho=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon$. Furthermore, the bound is tight.
@@
The $L_{p}$ ($p=1,2$) sensitivity of a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is given by $\Delta_{p}(f)=\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}}\|f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})-f(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\|_{p}$, where $\|\mathbf{v}\|_{p}$ is the $L_{p}$ norm of the vector $\mathbf{v}$.
Theorem 2.6 (Gaussian Mechanism of zCDP [8]).
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be a function of the dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, the Gaussian mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ adds i.i.d. Gaussian noise to $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ with scale parameter $\sigma^{2}$, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})+N(0,\sigma^{2}\mathbf{I}_{n}),$ where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. The Gaussian mechanism satisfies $\frac{\Delta_{2}(f)^{2}}{2\sigma^{2}}$-zCDP.